### v1

# Orthogonal projectors of binary LCD codes

Abstract
We prove that binary even LCD code and some graphs are in one-to-one correspondence in a certain way. Furthermore, we show that adjacency matrices of non-isomorphic simple graphs give inequivalent binary LCD codes, and vice versa.
1 Introduction
The interplay between codes and designs have provided many useful results [Assmus1992-rb]. Concerning relations between codes and graphs, several results are known: The rank of an adjacency matrix $A$ over $\mathbb{F}_{p}$ is called the $p$-rank of $A$. Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of adjacency matrices of strongly regular graphs are characterized by their eigenvalues. Haemers, Peeters and Rijckevorsel [Haemers1999-no] studied binary codes from adjacency matrices of strongly regular graphs. For strongly regular graphs with certain parameters, they observed that non-isomorphic graphs give inequivalent codes. Massey [massey_linear_1992] introduced linear complementary dual codes (LCD codes for short), and showed that a linear code has the orthogonal projector if and only if it is an LCD code. Since any orthogonal projector is symmetric matrix, adjacency matrices of some undirected graphs should coincide with orthogonal projectors. Based on this observation, Rodrigues and Keys studied a relation between LCD codes and adjacency matrices of graphs. They gave two conditions where adjacency matrices of graphs generate LCD codes over $\mathbb{F}_{q}$, where $q$ is a prime power [Key2018-ec, Proposition 2].
In this paper, we consider binary LCD codes in particular, and study a relation between binary LCD codes and adjacency matrices of simple graphs. It is known that the orthogonal projector $A$ of a binary LCD code is a symmetric matrix with $A^{2}=A$, where we regard $A$ as a matrix over $\mathbb{F}_{2}$. Hence any adjacency matrix satisfies the latter condition is an orthogonal projector of binary LCD code, as shown in [Key2018-ec, Brouwer1992-nc]. We show that such adjacency matrix is the orthogonal projector of a binary even LCD code, and two binary even LCD codes are equivalent if and only if the corresponding graphs are isomorphic. This establishes a one-to-one correspondence between binary even LCD codes and the above-mentioned class of graphs. Furthermore, we give a lower bound on any linear code obtained as a row span of strongly regular graphs. As an application, we consider codes generated by $srg(41,20,9,10)$, and partially solve a conjecture by Haemers, Peeters and Rijckevorsel [Haemers1999-no]. Furthermore, we improve some lower bounds on the minimum weight of binary LCD codes by constructing optimal binary $[n,k,d]$ LCD codes with the following parameters:
$\begin{split}(n,k,d)=&(31,10,10),(31,11,10),(31,20,6),(33,10,12),\\ &(33,11,11),(33,12,10),(33,13,10),(33,20,6),\\ &(33,22,6),(33,21,6),(34,13,10),(34,22,6).\end{split}$
We note that orthogonal projectors of the above optimal codes are adjacency matrices of regular graphs. By the above example, we show that from optimal binary LCD codes we can generate graphs whose adjacency matrices give codes with large minimum weight.
This paper is organized as follows: In Section 2, we introduce definitions and basic results of LCD codes and strongly regular graphs. In Section 3, we prove the main results. In Section 4, we apply the main results to a study of linear codes obtained as row spans of adjacency matrices of strongly regular graphs, and to a construction of optimal binary LCD codes.
2 Preliminaries
Let $\mathbb{F}_{q}$ be the finite field of order $q$, where $q$ is a prime power. Let $\mathbb{F}_{q}^{n}$ be the vector space of all $n$-tuples over $\mathbb{F}_{q}$. A $k$-dimensional subspace of $\mathbb{F}_{q}^{n}$ is said to be an $[n,k]$ code over $\mathbb{F}_{q}$. Especially, codes over $\mathbb{F}_{2}$ are said to be binary codes. Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$. The parameters $n$ and $k$ are said to be the length and the dimension of $C$ respectively. A vector in $C$ is called a codeword. The weight of $x=(x_{1},x_{2},\dots,x_{n})\in\mathbb{F}_{q}^{n}$ is defined as $\operatorname{wt}(x)=|\{i\mid x_{i}\neq 0\}|$. The minimum weight $d(C)$ of $C$ is the minimum weight among all nonzero codewords in $C$. If the minimum weight of $C$ equals to $d$, then $C$ is said to be an $[n,k,d]$ code over $\mathbb{F}_{q}$. We say that a code $C$ is even if all codewords have even weights; otherwise, the code $C$ is odd. Two $[n,k]$ codes $C_{1}$ and $C_{2}$ over $\mathbb{F}_{q}$ are equivalent if there is a monomial matrix $M$ such that $C_{2}=\{cM\mid c\in C_{1}\}$. The equivalence of two codes $C_{1}$ and $C_{2}$ is denoted by $C_{1}\simeq C_{2}$. A generator matrix of a code $C$ is any matrix whose rows form a basis of $C$.
The dual code $C^{\perp}$ of an $[n,k]$ code $C$ over $\mathbb{F}_{q}$ is defined as $C^{\perp}=\{x\in\mathbb{F}_{q}^{n}\mid(x,y)=0\text{ for all }y\in C\}$, where $(x,y)$ denotes the standard inner product. An $[n,k]$ code $C$ over $\mathbb{F}_{q}$ is said to be an linear complementary dual code (LCD code for short) if $C\cap C^{\perp}=\{\mathbf{0}_{n}\}$. Massey [massey_linear_1992] introduced the concept of LCD codes, and gave the following characterization:
Theorem 2.1 (Proposition 1 [massey_linear_1992]).
Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$ and let $G$ be a generator matrix of $C$. Then $C$ is an LCD code if and only if the $k\times k$ matrix $GG^{T}$ is nonsingular. Moreover, if $C$ is an LCD code, then $\Pi_{c}=G^{T}(GG^{T})^{-1}G$ is the orthogonal projector from $\mathbb{F}_{q}^{n}$ to $C$.
Let $C$ be a code over $\mathbb{F}_{q}$ with length $n$ and let $T\subseteq\{1,2,\dots,n\}$. Deleting every coordinate $i\in T$ in every codeword of $C$ gives a code called the punctured code on $T$ and denoted by $C^{T}$. Define a subcode $C(T)=\{(c_{1},c_{2},\dots,c_{n})\in C\mid c_{i}=0\text{ for all }i\in T\}$. Puncturing $C(T)$ on $T$ gives a code called the shortened code on $T$ and denoted by $C_{T}$. If $|T|=1$, say $T=\{i\}$, then we will write $C^{\{i\}}$ and $C_{\{i\}}$ as $C^{i}$ and $C_{i}$, respectively. Bouyuklieva [bouyuklieva_optimal_2021, Proposition 1] showed that any binary LCD $[n,k,d]$ code can be extended to binary even LCD $[n+1,k,d+1]$ code provided $k$ is even. This result is based on the characterization by Carlet, Mesnager, Tang and Qi [carlet_new_2019, Theorem 3]. Furthermore, Bouyuklieva gave a characterization of both punctured codes and shortened codes of binary LCD codes:
Theorem 2.2 (Lemma 2 [bouyuklieva_optimal_2021]).
Let $C$ be a binary LCD $[n,k]$ code with $d(C),d(C^{\perp})\geq 2$. For all $1\leq i\leq n$, exactly one of $C^{i}$ and $C_{i}$ is an LCD code.
Theorem 2.3 (Proposition 2 [bouyuklieva_optimal_2021]).
Let $C$ be a binary even LCD $[n,k]$ code with $d(C),d(C^{\perp})\geq 2$. Then the punctured code of C on any coordinate is again an LCD code.
A graph $\Gamma=(V,E)$ consists of a finite set of vertices $V$ together with a set of edges $E\subset\binom{V}{2}$. In this paper, we are concerned only with simple graphs, namely, graphs which are undirected, and without loops or multiple edges. We say that $x,y\in V$ are adjacent if $\{x,y\}\in E$, and denote by $x\sim y$. A vertex $y\in V$ is a neighbor of $x\in V$ if $x$ and $y$ are adjacent. The cardinality of the set of neighbors $N(x)=\{y\in V\mid x\sim y\}$ is called the valency of $x\in V$. A graph is $k$-regular if all its vertices have the same valency $k$. A simple regular graph is strongly regular with parameters $(v,k,\lambda,\mu)$ if it has $v$ vertices, valency $k$, and if any two adjacent vertices are together adjacent to $\lambda$ vertices, while any two nonadjacent vertices are together adjacent to $\mu$ vertices. A strongly regular graph with parameters $(v,k,\lambda,\mu)$ is denoted by $srg(v,k,\lambda,\mu)$. An adjacency matrix $A$ of a graph $\Gamma=(V,E)$ is a $|V|\times|V|$ matrix with rows and columns labelled by the vertices $x,y\in V$, defined by $A_{x,y}=1$ if $x\sim y$ and $A_{x,y}=0$ otherwise. Two graphs $\Gamma$ and $\Gamma^{\prime}$ with adjacency matrices $A$ and $A^{\prime}$ are equivalent if there is a permutation matrix $P$ such that $A^{\prime}=P^{T}AP$.
The rank of an adjacency matrix $A$ over $\mathbb{F}_{p}$ is called the $p$-rank of $A$. Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of adjacency matrices of strongly regular graphs are characterized by their eigenvalues. Haemers, Peeters and Rijckevorsel [Haemers1999-no] considered binary codes obtained as row spans of strongly regular graphs, and obtained many computational results for strongly regular graphs with small vertices. Keys and Rodrigues [Key2018-ec] constructed many LCD codes over $\mathbb{F}_{q}$ from adjacency matrices of strongly regular graphs via orthogonal projectors. Especially, they introduced the concept of reflexive LCD codes and complementary LCD codes [Key2018-ec, Definition 2], and provided the condition where adjacency matrices of graphs give LCD codes over $\mathbb{F}_{q}$ [Key2018-ec, Proposition 2].
3 Main results
In this section, we give main results on a correspondance between binary even LCD codes and adjacency matrices of certain simple graphs. Also, we provide a lower bound on the minimum weights of binary codes that come from strongly regular graphs.
3.1 Characterization
Lemma 3.1.
Let $C$ be a binary LCD code with orthogonal projector $A$. Then $C$ is the row span of $A$.
Proof.
Let $e_{i}$ be the unit vector whose componets are all zero except $i$-th component, and let $r_{i}$ be the $i$-th row of $A$. Then, $Ae_{i}=r_{i}\in C$ for all $1\leq i\leq n$. Hence the row span of $A$ is a subspace of $C$. On the other hand, for any $c\in C$ it holds that $Ac=c$. This tells us that any codeword $c$ can be expressed as a linear combination of the rows of $A$. Hence $C$ is a subspace of the row span of $A$. This completes the proof. ∎
Lemma 3.2.
Let $C$ and $C^{\prime}$ be binary LCD codes with orthogonal projectors $A$ and $A^{\prime}$, respectively. Then, $C$ is equivalent to $C^{\prime}$ if and only if there is a permutation matrix $P$ such that $A^{\prime}=P^{T}AP$.
Proof.
Let $G$ and $G^{\prime}$ be generator matrices of $C$ and $C^{\prime}$ respectively. If $C$ and $C^{\prime}$ are equivalent, then there are a monomial matrix $M$ and a permutation matrix $P$ such that $G^{\prime}=MGP$. By Theorem 2.1, $A^{\prime}=P^{T}AP$ follows.
Conversely, suppose that there is a permutation matrix $P$ such that $A^{\prime}=P^{T}AP$. Any permutation of the rows of $A$, $P^{T}$ in particular, preserves its row span. By Lemma 3.1, the row spans of $A$ and $A^{\prime}$ are $C$ and $C^{\prime}$ respectively. It follows that $C^{\prime}=CP$, and hence $C$ and $C^{\prime}$ are equivalent, as required. ∎
Corollary 3.3.
Let $C$ and $C^{\prime}$ be binary LCD codes with orthogonal projectors $A$ and $A^{\prime}$, respectively. If the traces of $A$ and $A^{\prime}$ are distinct, then $C$ and $C^{\prime}$ are inequivalent.
Remark 3.4.
By the proof of Lemma 3.2, it follows that the orthogonal projector of an LCD code is independent of the choice of a generator matrix.
Remark 3.5.
Any adjacency matrix $A$ of a simple graph satisfying $A^{2}=A$ is the orthogonal projector of a binary LCD code, where we regard $A$ as a matrix over $\mathbb{F}_{2}$. These graphs are characterized by the following conditions:
  • If two vertices $x$ and $y$ are adjacent, then the number of the commmon neighbors of $x$ and $y$ is odd.
  • If two vertices $x$ and $y$ are not adjacent, then the number of the commmon neighbors of $x$ and $y$ is even.
Among these graphs there are strongly regular graphs with parameters $(v,k,\lambda,\mu)$, where $k,\mu\equiv 0\pmod{2}$ and $\lambda\equiv 1\pmod{2}$.
Lemma 3.6.
Let $C$ be a binary LCD code with orthogonal projector $A$. Then $A$ is an adjacency matrix of a simple graph if and only if $C$ is an even code.
Proof.
Let $r_{i}$ be the $i$-th row of $A$, and let $n$ be the length of $C$. If $A$ is an adjacency matrix of a simple graph, then any diagonal entry of $A$ is zero. This fact, along with $A^{2}=A$, implies $(r_{i},r_{i})=0$ for all $1\leq i\leq n$. That is, any row of $A$ has an even weight. By Lemma 3.1, $C$ is even.
Conversely, suppose $C$ is even. Since the orthogonal projector $A$ is a symmetric $(0,1)$-matrix, it remains to show that any diagonal entry of $A$ is 00. This can be done by the above argument along with the fact that every codeword of $C$ has an even weight and $A^{2}=A$. This completes the proof. ∎
Theorem 3.7.
The orthogonal projector $A$ of a binary even LCD code is an adjacency matrix of a simple graph satisfying $A^{2}=A$, where we regard $A$ as a matrix over $\mathbb{F}_{2}$. Conversely, such an adjacency matrix is the orthogonal projector of a binary even LCD code. Furthermore, two binary even LCD codes are equivalent if and only if corresponding graphs are isomorphic.
Proof.
This immediately follows from Lemmas 3.1, 3.2 and 3.6. ∎
Remark 3.8.
Bouyuklieva studied punctured codes and shortened codes of binary LCD codes and obtained Theorem 2.2. To prove the theorem, they essentially considered the orthogonal projection of the unit vector $e_{i}$ and made the following observation: If the $i$-th component of the orthogonal projection of $e_{i}$ is $1$, then $C_{i}$ is LCD and $C^{i}$ is not LCD; otherwise, $C^{i}$ is LCD and $C_{i}$ is not LCD. The $i$-th component of the orthogonal projection of $e_{i}$ is equal to the $i$-th diagonal entry of the orthogonal projector $A$. Hence it follows that $(A)_{i,i}=0$ if and only if $C^{i}$ is LCD, and $A_{i,i}=1$ if only if $C_{i}$ is LCD.
3.2 Bounds On Minimum Weights
In this section, we give a lower bound on the minimum weights of binary codes obtained as row spans of adjacency matrices of strongly regular graphs. These results will be applied in the following sections. Note that all the assertions in this section hold not only for binary LCD codes but also for any binary codes.
Proposition 3.9.
Let $A$ be an adjacency matrix of $srg(v,k,\lambda,\mu)$, and let $C$ be a code with generator matrix $A$. Then $d(C^{\perp})\geq 1+k/t$, where $t=\max\{\lambda,\mu\}$.
Proof.
If $d=d(C^{\perp})$, then there is a set of $d$ linearly dependent columns of $A$. Let $S$ be a submatrix of $A$ consists of such columns, and $n_{i}$ be the number of rows in $S$ with weight $i$. Since the sum of components of every row of $S$ equals to zero, the weight of every row of $S$ must be even. Counting in two ways the number of choices of $j$ 1s in each row of $S$ for $j=1,2$, we have
$\begin{split}\sum_{i}(2i)n_{2i}&=dk,\\ \sum_{i}(2i)(2i-1)n_{2i}&\leq d(d-1)t.\\ \end{split}$
It follows that $0\leq\sum_{i}2i(2i-2)\leq d((d-1)m-k)$, and hence $d\geq 1+k/m$. ∎
Proposition 3.10.
Let $A$ be an adjacency matrix of $srg(v,k,\lambda,\mu)$, and let $C$ be a code with generator matrix $A+I$. Then $d(C^{\perp})\geq 1+(k+1)/(t+1)$, where $t=\max\{\lambda,\mu\}$.
Proof.
With the notation of that in the proof of Proposition 3.9, count in two ways the number of choices of $j$ 1s in each row of $S$ for $j=1,2$. Then we have
$\begin{split}\sum_{i}(2i)n_{2i}&=d(k+1),\\ \sum_{i}(2i)(2i-1)n_{2i}&\leq d(d-1)(t+1).\\ \end{split}$
It follows that $d\geq 1+(k+1)/(t+1)$. ∎
Corollary 3.11.
Let $A$ be an adjacency matrix of $srg(v,k,\lambda,\mu)$, and let $C$ be a code with generator matrix $A$. If $k$ is odd, then $d(C^{\perp})$ is even.
Proof.
With the notation of that in the proof of Proposition 3.9, suppose that $d$ is odd. Counting in two ways the number of choices of 1s in each row of $S$, we have
$\sum_{i}(2i)n_{2i}=dk.$
However, this is absurd since the left hand side of the equation is even, while the right hand side is odd. This completes the proof. ∎
4 Applications
In this section, we illustrate examples where results from the preceding section can be applied.
4.1 Binary Codes from $srg(41,20,9,10)$
Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of strongly regular graphs are characterized by their eigenvalues. For strongly regular graphs $srg(41,20,9,10)$, any corresponding binary code has $2$-rank, namely, dimension $20$. Haemers, Peeters and Rijckevorsel [Haemers1999-no] considered binary codes obtained as row spans of strongly regular graphs. They observed that adjacency matrices of known non-isomorphic $srg(41,20,9,10)$ gave inequivalent codes, and among such codes the code from the Paley graph had the largest minimum weight. They conjectured that the latter statement holds in general.
Remarkably, the former statement holds in general for any strongly regular graph with parameters $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ for $q\equiv 1\pmod{8}$ since any adjacency matrix of such a graph satisfies the condition given in Theorem 3.7. Furthermore, we can show that the Paley graph gives a binary even $[40,20]$ code by the following argument: The Griesmer bound shows that the largest minimum weight of binary LCD $[41,20]$ code is $11$. By Lemma 3.6, the minimum weight must be even. This proves that the Paley graph indeed gives a code with largest minimum weight among all $srg(41,20,9,10)$. Note that the classification of $srg(41,20,9,10)$ has not been completed yet. We state this fact as the following theorem:
Theorem 4.1.
Any pair of non-isomorphic $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ with $q\equiv 1\pmod{8}$ gives a pair of inequivalent binary even LCD codes, and vice versa. Furthermore, the Paley graph of order $41$ gives a binary even $[40,20,10]$ code whose minimum weight is largest among all $srg(40,20,9,10)$.
According to [Maksimovic2022-xa], there are at least $7152$ strongly regular graphs with parameters $(41,20,9,10)$ up to isomorphism. We constructed binary even LCD codes from the $7152$ graphs, and verified the following:
  • There are mutually inequivalent,
  • Only the code that comes from the Paley graph has the largest minimum weight $10$,
  • All codes have minimum weight greater than $4$, and the code comes from the following adjacency matrix has the minimum weight $4$.
By Proposition 3.10, any code from $srg(41,20,9,10)$ must have the minimum weight larger than $d\geq 1+(k+1)/(\min\{\lambda,\mu\}+1)=3$. Applying Theorem 3.7, we can improve the lower bound by $1$, and hence the lower bound on the minimum weight is $4$. The above observation shows this lower bound is tight for $srg(40,20,9,10)$.
4.2 Construction by Orthogonal Projectors
In this section, we improve some lower bounds on the minimum weight of binary LCD codes by constructing optimal binary LCD codes from their orthogonal projectors. For recent results on optimal binary LCD codes, we refer the reader to [bouyuklieva_optimal_2021]. Let $d(n,k)$ denote the largest minimum weight among all $[n,k]$ LCD code.
Here we consider a construction by circulant matrices. A circulant matrix $C$ of order $n$ is a matrix of the following form
$C=\begin{pmatrix}c_{0}&c_{1}&c_{2}&\cdots&c_{n-2}&c_{n-1}\\ c_{n-1}&c_{0}&c_{1}&\cdots&c_{n-3}&c_{n-2}\\ c_{n-2}&c_{n-1}&c_{0}&\cdots&c_{n-4}&c_{n-3}\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ c_{2}&c_{3}&c_{4}&\cdots&c_{0}&c_{1}\\ c_{1}&c_{2}&c_{3}&\cdots&c_{n-1}&c_{0}\\ \end{pmatrix}.$
Circulant matrices are determined by their first rows, and hence there are $2^{n}$ circulant matrices over $\mathbb{F}_{2}$ of order $n$. For $26\leq n\leq 37$, we constructed all circulant matrices over $\mathbb{F}_{2}$ of order $n$, which must be checked further for the condition given in Theorem 3.7. Then by Lemma 3.1 we constructed the binary LCD codes from the orthogonal projectors, and computed their minimum weights. Consequently, we obtain the following theorem:
Theorem 4.2.
There are optimal binary LCD $[n,k,d]$ codes for
$(n,k,d)=$ $(31,10,10),(31,11,10),(31,20,6),(33,10,12),$
$(33,11,11),(33,12,10),(33,13,10),(33,20,6),(33,22,6).$
For each of these codes, the first rows of their orthogonal projectors are listed in Table 1. For a binary $[n,k,d]$ LCD code $C$, it is easy to see that $C^{\prime}=\{(0,c)\mid c\in C\}$ is a binary LCD $[n+1,k,d]$ code. Using this property, we obtained three optimal binary LCD codes from some codes listed in Theorem 4.2. Consequently, we obtain the following theorem:
Theorem 4.3.
There are optimal binary LCD $[n,k,d]$ codes for
$(n,k,d)=(33,21,6),(34,13,10),(34,22,6).$
Any orthogonal projector mentioned above has constant diagonal entries, and hence any binary LCD code in Theorem 4.2 is either an even code or its dual code. Note that the orthogonal projectors of LCD codes above (or their dual codes) are adjacent matrices of regular graphs by construction.
Table 1: Optimal binary even LCD codes
$n$ $k$ $d$ the first row
$31$ $10$ $10$ $(0000010001110100001011100010000)$
$31$ $11$ $10$ $(1001011101111110011111101110100)$
$31$ $20$ $6$ $(0001011101111110011111101110100)$
$33$ $10$ $12$ $(000101110111111100111111101110100)$
$33$ $11$ $11$ $(100001010010011000011001001010000)$
$33$ $12$ $10$ $(000101110110111100111101101110100)$
$33$ $13$ $10$ $(100001010011011000011011001010000)$
$33$ $20$ $6$ $(000001010011011000011011001010000)$
$33$ $22$ $6$ $(000001010010011000011001001010000)$
5 Concluding Remarks
In this paper, we showed that binary even LCD codes and certain graphs are in one-to-one correspondence in a certain sense. Furthermore, we provided lower bounds on minimum weights of codes obtained from strongly regular graphs. Consequently, a part of the conjecture by Haemers, Peeters, Rijckevorsel [Haemers1999-no] was solved. The conjecture that the Paley graph gives a code with the largest weight among all the strongly regular graphs is still open for $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ with $q\equiv 1\pmod{8}$ and $q>41$. Our result can be applied to generate a graph that gives optimal binary code, as shown in Section 4. We hope that studying such graphs leads to a discovery of a further relation between codes and graphs.

### v2

@@
 Furthermore, we show that adjacency matrices of non-isomorphic simple graphs give inequivalent binary LCD codes, and vice versa.
+Keywords: LCD code, strongly regular graph.
+Mathematical subject classification (2020): 94B05, 05C50.
 1 Introduction
@@
 Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of adjacency matrices of strongly regular graphs are characterized by their eigenvalues.
-Haemers, Peeters and Rijckevorsel [Haemers1999-no] studied binary codes from adjacency matrices of strongly regular graphs.
-For strongly regular graphs with certain parameters, they observed that non-isomorphic graphs give inequivalent codes.
 Massey [massey_linear_1992] introduced linear complementary dual codes (LCD codes for short), and showed that a linear code has the orthogonal projector if and only if it is an LCD code.
@@
 This establishes a one-to-one correspondence between binary even LCD codes and the above-mentioned class of graphs.
-Furthermore, we give a lower bound on any linear code obtained as a row span of strongly regular graphs.
-As an application, we consider codes generated by $srg(41,20,9,10)$, and partially solve a conjecture by Haemers, Peeters and Rijckevorsel [Haemers1999-no].
-Furthermore, we improve some lower bounds on the minimum weight of binary LCD codes by constructing optimal binary $[n,k,d]$ LCD codes with the following parameters:
+Haemers, Peeters and Rijckevorsel [Haemers1999-no] studied binary codes from adjacency matrices of strongly regular graphs.
+For strongly regular graphs with certain parameters, they observed that non-isomorphic graphs give inequivalent codes.
+This paper gives an explanation to the above observation.
+Furthermore, we improve some lower bounds on the minimum weight of binary LCD codes by constructing optimal binary LCD $[n,k,d]$ codes with the following parameters:
 $\begin{split}(n,k,d)=&(31,10,10),(31,11,10),(31,20,6),(33,10,12),\\ &(33,11,11),(33,12,10),(33,13,10),(33,20,6),\\ &(33,22,6),(33,21,6),(34,13,10),(34,22,6).\end{split}$
 We note that orthogonal projectors of the above optimal codes are adjacency matrices of regular graphs.
-By the above example, we show that from optimal binary LCD codes we can generate graphs whose adjacency matrices give codes with large minimum weight.
+By the above example, we show that from optimal binary even LCD codes we can generate graphs whose adjacency matrices give codes with large minimum weight.
 This paper is organized as follows: In Section 2, we introduce definitions and basic results of LCD codes and strongly regular graphs.
@@
 Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$.
-The parameters $n$ and $k$ are said to be the length and the dimension of $C$ respectively.
+The parameters $n$ and $k$ are said to be the length and the dimension of $C$, respectively.
 A vector in $C$ is called a codeword.
@@
 Puncturing $C(T)$ on $T$ gives a code called the shortened code on $T$ and denoted by $C_{T}$.
-If $|T|=1$, say $T=\{i\}$, then we will write $C^{\{i\}}$ and $C_{\{i\}}$ as $C^{i}$ and $C_{i}$, respectively.
-Bouyuklieva [bouyuklieva_optimal_2021, Proposition 1] showed that any binary LCD $[n,k,d]$ code can be extended to binary even LCD $[n+1,k,d+1]$ code provided $k$ is even.
-This result is based on the characterization by Carlet, Mesnager, Tang and Qi [carlet_new_2019, Theorem 3].
-Furthermore, Bouyuklieva gave a characterization of both punctured codes and shortened codes of binary LCD codes:
+If $|T|=1$, say $T=\{i\}$, then we will write $C^{\{i\}}$ and $C_{\{i\}}$ simply as $C^{i}$ and $C_{i}$, respectively.
+Bouyuklieva gave a characterization of both punctured codes and shortened codes of binary LCD codes:
 Theorem 2.2 (Lemma 2 [bouyuklieva_optimal_2021]).
@@
 The rank of an adjacency matrix $A$ over $\mathbb{F}_{p}$ is called the $p$-rank of $A$.
-Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of adjacency matrices of strongly regular graphs are characterized by their eigenvalues.
-Haemers, Peeters and Rijckevorsel [Haemers1999-no] considered binary codes obtained as row spans of strongly regular graphs, and obtained many computational results for strongly regular graphs with small vertices.
+Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of adjacency matrices of some strongly regular graphs are characterized by their eigenvalues.
+Haemers, Peeters and Rijckevorsel [Haemers1999-no] considered binary codes obtained as row spans of strongly regular graphs, and obtained several computational results for strongly regular graphs with small vertices.
 Keys and Rodrigues [Key2018-ec] constructed many LCD codes over $\mathbb{F}_{q}$ from adjacency matrices of strongly regular graphs via orthogonal projectors.
@@
 3 Main results
-In this section, we give main results on a correspondance between binary even LCD codes and adjacency matrices of certain simple graphs.
+In this section, we give main results on a correspondence between binary even LCD codes and adjacency matrices of some simple graphs.
 Also, we provide a lower bound on the minimum weights of binary codes that come from strongly regular graphs.
@@
 Proof.
-Let $e_{i}$ be the unit vector whose componets are all zero except $i$-th component, and let $r_{i}$ be the $i$-th row of $A$.
+Let $e_{i}$ be the unit vector whose components are all zero except $i$-th component, and let $r_{i}$ be the $i$-th row of $A$.
 Then, $Ae_{i}=r_{i}\in C$ for all $1\leq i\leq n$.
@@
 These graphs are characterized by the following conditions:
-  • If two vertices $x$ and $y$ are adjacent, then the number of the commmon neighbors of $x$ and $y$ is odd.
-  • If two vertices $x$ and $y$ are not adjacent, then the number of the commmon neighbors of $x$ and $y$ is even.
+  • If two vertices $x$ and $y$ are adjacent, then the number of the common neighbors of $x$ and $y$ is odd.
+  • If two vertices $x$ and $y$ are not adjacent, then the number of the common neighbors of $x$ and $y$ is even.
 Among these graphs there are strongly regular graphs with parameters $(v,k,\lambda,\mu)$, where $k,\mu\equiv 0\pmod{2}$ and $\lambda\equiv 1\pmod{2}$.
@@
 Remark 3.8.
-Bouyuklieva studied punctured codes and shortened codes of binary LCD codes and obtained Theorem 2.2.
+Bouyuklieva studied punctured codes and shortened codes of binary LCD codes and obtained Theorem 2.2 (see Section 2).
 To prove the theorem, they essentially considered the orthogonal projection of the unit vector $e_{i}$ and made the following observation: If the $i$-th component of the orthogonal projection of $e_{i}$ is $1$, then $C_{i}$ is LCD and $C^{i}$ is not LCD; otherwise, $C^{i}$ is LCD and $C_{i}$ is not LCD.
@@
 3.2 Bounds on minimum weights
-In this section, we give a lower bound on the minimum weights of binary codes obtained as row spans of adjacency matrices of strongly regular graphs.
-These results will be applied in the following sections.
-Note that all the assertions in this section hold not only for binary LCD codes but also for any binary codes.
+In this subsection, we give a lower bound on the minimum weights of binary codes obtained as row spans of adjacency matrices of strongly regular graphs.
+These results will be applied in Section 4.
+Note that all the assertions in this subsection hold not only for binary LCD codes but also for any binary codes.
 Proposition 3.9.
@@
 $\sum_{i}(2i)n_{2i}=dk.$
-However, this is absurd since the left hand side of the equation is even, while the right hand side is odd. This completes the proof. ∎
+However, this is absurd since the left hand side of the equation is even while the right hand side is odd. This completes the proof. ∎
 4 Applications
 In this section, we illustrate examples where results from the preceding section can be applied.
-4.1 Binary Codes from $srg(41,20,9,10)$
-Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of strongly regular graphs are characterized by their eigenvalues.
+4.1 Binary codes from the Paley graph
+Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of adjacency matrices of strongly regular graphs are characterized by their eigenvalues.
 For strongly regular graphs $srg(41,20,9,10)$, any corresponding binary code has $2$-rank, namely, dimension $20$.
 Haemers, Peeters and Rijckevorsel [Haemers1999-no] considered binary codes obtained as row spans of strongly regular graphs.
-They observed that adjacency matrices of known non-isomorphic $srg(41,20,9,10)$ gave inequivalent codes, and among such codes the code from the Paley graph had the largest minimum weight.
-They conjectured that the latter statement holds in general.
-Remarkably, the former statement holds in general for any strongly regular graph with parameters $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ for $q\equiv 1\pmod{8}$ since any adjacency matrix of such a graph satisfies the condition given in Theorem 3.7.
-Furthermore, we can show that the Paley graph gives a binary even $[40,20]$ code by the following argument: The Griesmer bound shows that the largest minimum weight of binary LCD $[41,20]$ code is $11$.
-By Lemma 3.6, the minimum weight must be even.
-This proves that the Paley graph indeed gives a code with largest minimum weight among all $srg(41,20,9,10)$.
-Note that the classification of $srg(41,20,9,10)$ has not been completed yet.
-We state this fact as the following theorem:
-Theorem 4.1.
+They observed that for $srg(25,12,5,6)$ and $srg(41,20,9,10)$, any pair of non-isomorphic graphs gave a pair of inequivalent codes, and among such codes the code from the Paley graph had the largest minimum weight.
+Based on this result, they conjectured that the latter statement holds in general (see [Haemers1999-no, Section 4.3]).
+The classification of $srg(25,12,5,6)$ has been completed [paulus1973conference]; there are exactly $15$ non-isomorphic $srg(25,12,5,6)$, and with the aid of computer it was shown in [Haemers1999-no, Section 6.1] that the Paley graph gives the code with largest minimum weights.
+On the other hand, the classification of $srg(41,20,9,10)$ has not been completed yet.
+However, it is not difficult to show the Paley graph of order $41$ gives a binary even $[41,20]$ code whose minimum weight is largest among all $srg(41,20,9,10)$.
+One way to see this is to use the Griesmer bound, which tells us the minimum weight of binary LCD $[41,20]$ code is at most $11$.
+By Lemma 3.6, the minimum weight must be even, and hence the result follows.
+According to [grassl_bounds_nodate], the largest minimum weight of binary linear code is exactly $10$, so it is obvious that the code from the Paley graph of order $41$ attains the largest minimum weight.
+To best of our knowledge, the conjecture is still open for $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ with $q>41$ and $q\equiv 1\pmod{8}$.
+Remarkably, the first part of the above observation by Haemers, Peeters and Rijckevorsel holds in general for any strongly regular graph with parameters $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ for $q\equiv 1\pmod{8}$ since any adjacency matrix of such a graph satisfies the condition given in Theorem 3.7.
+We state this fact as the following corollary:
+Corollary 4.1.
 Any pair of non-isomorphic $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ with $q\equiv 1\pmod{8}$ gives a pair of inequivalent binary even LCD codes, and vice versa.
-Furthermore, the Paley graph of order $41$ gives a binary even $[40,20,10]$ code whose minimum weight is largest among all $srg(40,20,9,10)$.
 According to [Maksimovic2022-xa], there are at least $7152$ strongly regular graphs with parameters $(41,20,9,10)$ up to isomorphism.
@@
 Applying Theorem 3.7, we can improve the lower bound by $1$, and hence the lower bound on the minimum weight is $4$.
-The above observation shows this lower bound is tight for $srg(40,20,9,10)$.
-4.2 Construction by Orthogonal Projectors
-In this section, we improve some lower bounds on the minimum weight of binary LCD codes by constructing optimal binary LCD codes from their orthogonal projectors.
+The above observation shows this lower bound is tight for $srg(41,20,9,10)$.
+4.2 Optimal LCD codes from regular graphs
+In this subsection, we improve some lower bounds on the minimum weight of binary LCD codes by constructing optimal binary LCD codes from their orthogonal projectors.
 For recent results on optimal binary LCD codes, we refer the reader to [bouyuklieva_optimal_2021].
-Let $d(n,k)$ denote the largest minimum weight among all $[n,k]$ LCD code.
+Let $d(n,k)$ denote the largest minimum weight among all LCD $[n,k]$ code.
 Here we consider a construction by circulant matrices. A circulant matrix $C$ of order $n$ is a matrix of the following form
@@
 For each of these codes, the first rows of their orthogonal projectors are listed in Table 1.
-For a binary $[n,k,d]$ LCD code $C$, it is easy to see that $C^{\prime}=\{(0,c)\mid c\in C\}$ is a binary LCD $[n+1,k,d]$ code.
+For a binary LCD $[n,k,d]$ code $C$, it is easy to see that $C^{\prime}=\{(0,c)\mid c\in C\}$ is a binary LCD $[n+1,k,d]$ code.
 Using this property, we obtained three optimal binary LCD codes from some codes listed in Theorem 4.2.
@@
 Note that the orthogonal projectors of LCD codes above (or their dual codes) are adjacent matrices of regular graphs by construction.
+It is remarkable that adjacency matrices of regular graphs can give LCD codes with the largest minimum weights.
+Furthermore, we remark that any code in Theorem 4.2 has odd length.
+This follows by the fact that the dimension of any binary even LCD code is even, and that the dual code of any even code must contain the all-ones vector; if a binary even LCD $[n,k]$ code $C$ has even length, then the all-ones vector must be in $C\cap C^{\perp}$, a contradiction.
 Table 1: Optimal binary even LCD codes
@@
 5 Concluding remarks
-In this paper, we showed that binary even LCD codes and certain graphs are in one-to-one correspondence in a certain sense.
+In this paper, we proved that binary even LCD codes and some graphs are in one-to-one correspondence in a certain way.
 Furthermore, we provided lower bounds on minimum weights of codes obtained from strongly regular graphs.
-Consequently, a part of the conjecture by Haemers, Peeters, Rijckevorsel [Haemers1999-no] was solved.
-The conjecture that the Paley graph gives a code with the largest weight among all the strongly regular graphs is still open for $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ with $q\equiv 1\pmod{8}$ and $q>41$.
+Consequently, we gave an explanation to the observation by Haemers, Peeters, Rijckevorsel [Haemers1999-no].
 Our result can be applied to generate a graph that gives optimal binary code, as shown in Section 4.
-We hope that studying such graphs leads to a discovery of a further relation between codes and graphs.
+We hope that studying such graphs leads to a discovery of a further relation between graphs and even LCD codes.
+6 Acknowledgement
+The author would like to thank Stefka Bouyuklieva for several questions and comments that improved the manuscript.

### v3

@@
-Orthogonal projectors of binary LCD codes
+Orthogonal projectors of LCD codes and their graph representations
 Abstract
-We prove that binary even LCD code and some graphs are in one-to-one correspondence in a certain way.
-Furthermore, we show that adjacency matrices of non-isomorphic simple graphs give inequivalent binary LCD codes, and vice versa.
+We establish one-to-one correspondences between (i) binary even LCD codes and certain simple graphs, and (ii) ternary LCD codes and certain two-graphs.
 Keywords: LCD code, strongly regular graph.
@@
 1 Introduction
-The interplay between codes and designs have provided many useful results [Assmus1992-rb].
+The interplay between codes and designs has provided many useful results [Assmus1992-rb].
 Concerning relations between codes and graphs, several results are known: The rank of an adjacency matrix $A$ over $\mathbb{F}_{p}$ is called the $p$-rank of $A$.
@@
 Massey [massey_linear_1992] introduced linear complementary dual codes (LCD codes for short), and showed that a linear code has the orthogonal projector if and only if it is an LCD code.
-Since any orthogonal projector is symmetric matrix, adjacency matrices of some undirected graphs should coincide with orthogonal projectors.
-Based on this observation, Rodrigues and Keys studied a relation between LCD codes and adjacency matrices of graphs.
-They gave two conditions where adjacency matrices of graphs generate LCD codes over $\mathbb{F}_{q}$, where $q$ is a prime power [Key2018-ec, Proposition 2].
-In this paper, we consider binary LCD codes in particular, and study a relation between binary LCD codes and adjacency matrices of simple graphs.
+Since any orthogonal projector is a symmetric matrix, adjacency matrices of some undirected graphs should coincide with orthogonal projectors.
+Based on this observation, Rodrigues and Keys studied the relation between LCD codes and adjacency matrices of graphs.
+They gave two conditions under which adjacency matrices of graphs generate LCD codes over $\mathbb{F}_{q}$, where $q$ is a prime power [Key2018-ec, Proposition 2].
+In this paper, we establish one-to-one correspondences between binary even LCD codes and certain simple graphs; and ternary LCD codes and certain two-graphs.
 It is known that the orthogonal projector $A$ of a binary LCD code is a symmetric matrix with $A^{2}=A$, where we regard $A$ as a matrix over $\mathbb{F}_{2}$.
 Hence any adjacency matrix satisfies the latter condition is an orthogonal projector of binary LCD code, as shown in [Key2018-ec, Brouwer1992-nc].
-We show that such adjacency matrix is the orthogonal projector of a binary even LCD code, and two binary even LCD codes are equivalent if and only if the corresponding graphs are isomorphic.
-This establishes a one-to-one correspondence between binary even LCD codes and the above-mentioned class of graphs.
-Haemers, Peeters and Rijckevorsel [Haemers1999-no] studied binary codes from adjacency matrices of strongly regular graphs.
+We show that such an adjacency matrix is the orthogonal projector of a binary even LCD code, and two binary even LCD codes are equivalent if and only if the corresponding graphs are isomorphic.
+Furthermore, we prove an analogous results for the relation between ternary LCD codes and two-graphs.
+Haemers, Peeters, and Rijckevorsel [Haemers1999-no] studied binary codes from adjacency matrices of strongly regular graphs.
 For strongly regular graphs with certain parameters, they observed that non-isomorphic graphs give inequivalent codes.
 This paper gives an explanation to the above observation.
-Furthermore, we improve some lower bounds on the minimum weight of binary LCD codes by constructing optimal binary LCD $[n,k,d]$ codes with the following parameters:
-$\begin{split}(n,k,d)=&(31,10,10),(31,11,10),(31,20,6),(33,10,12),\\ &(33,11,11),(33,12,10),(33,13,10),(33,20,6),\\ &(33,22,6),(33,21,6),(34,13,10),(34,22,6).\end{split}$
-We note that orthogonal projectors of the above optimal codes are adjacency matrices of regular graphs.
-By the above example, we show that from optimal binary even LCD codes we can generate graphs whose adjacency matrices give codes with large minimum weight.
-This paper is organized as follows: In Section 2, we introduce definitions and basic results of LCD codes and strongly regular graphs.
-In Section 3, we prove the main results.
-In Section 4, we apply the main results to a study of linear codes obtained as row spans of adjacency matrices of strongly regular graphs, and to a construction of optimal binary LCD codes.
+The remaining part of this paper is organized as follows: In Section 2, we introduce definitions and basic results of LCD codes, simple graphs, and two-graphs.
+We give our main results in Section 3.
+Section 4 establishes auxiliary results on the minimum weight of codes derived from strongly regular graphs, which are later used in Section 5.
+In that section, we apply our main results to analyze binary LCD codes obtained as row spans of adjacency matrices of strongly regular graphs.
+Finally, Section 6 presents concluding remarks and open problems.
 2 Preliminaries
+2.1 LCD codes
 Let $\mathbb{F}_{q}$ be the finite field of order $q$, where $q$ is a prime power.
@@
 A $k$-dimensional subspace of $\mathbb{F}_{q}^{n}$ is said to be an $[n,k]$ code over $\mathbb{F}_{q}$.
-Especially, codes over $\mathbb{F}_{2}$ are said to be binary codes.
+In particular, codes over $\mathbb{F}_{2}$ and $\mathbb{F}_{3}$ are said to be binary and ternary codes, respectively.
 Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$.
@@
 The equivalence of two codes $C_{1}$ and $C_{2}$ is denoted by $C_{1}\simeq C_{2}$.
-A generator matrix of a code $C$ is any matrix whose rows form a basis of $C$.
+A generator matrix of a code $C$ is a matrix whose rows form a basis of $C$.
+The orthogonal projector $\Pi_{C}$ of an $[n,k]$ code $C$ over $\mathbb{F}_{q}$ is a unique linear map from $\mathbb{F}_{q}^{n}$ to $C$ satisfying
+$v\Pi_{C}=\begin{cases}v&\textrm{if }v\in C,\\ 0&\textrm{otherwise}.\end{cases}$
+An orthogonal projector is represented by a matrix $A$ satisfying $A^{T}=A$ and $A^{2}=A$.
+Conversely, any matrix satisfying $A^{T}=A$ and $A^{2}=A$ represents the orthogonal projector of some code.
+We refer the reader to [massey_linear_1992, bardet2019permutation] for more information on orthogonal projectors.
+In this paper, we often identify the orthogonal projector with its matrix representation in the standard basis.
 The dual code $C^{\perp}$ of an $[n,k]$ code $C$ over $\mathbb{F}_{q}$ is defined as $C^{\perp}=\{x\in\mathbb{F}_{q}^{n}\mid(x,y)=0\text{ for all }y\in C\}$, where $(x,y)$ denotes the standard inner product.
@@
 Puncturing $C(T)$ on $T$ gives a code called the shortened code on $T$ and denoted by $C_{T}$.
-If $|T|=1$, say $T=\{i\}$, then we will write $C^{\{i\}}$ and $C_{\{i\}}$ simply as $C^{i}$ and $C_{i}$, respectively.
-Bouyuklieva gave a characterization of both punctured codes and shortened codes of binary LCD codes:
+We will write $C^{\{i\}}$ and $C_{\{i\}}$ simply as $C^{i}$ and $C_{i}$, respectively.
+Bouyuklieva [bouyuklieva_optimal_2021] gave a characterization of both punctured codes and shortened codes of binary LCD codes:
 Theorem 2.2 (Lemma 2 [bouyuklieva_optimal_2021]).
@@
 Let $C$ be a binary even LCD $[n,k]$ code with $d(C),d(C^{\perp})\geq 2$. Then the punctured code of C on any coordinate is again an LCD code.
+2.2 Strongly regular graphs
+In this subsection, we review the definition and basic properties of strongly regular graphs.
+For a comprehensive treatment on this topic, we refer the reader to [cameron1991designs].
 A graph $\Gamma=(V,E)$ consists of a finite set of vertices $V$ together with a set of edges $E\subset\binom{V}{2}$.
@@
 A strongly regular graph with parameters $(v,k,\lambda,\mu)$ is denoted by $srg(v,k,\lambda,\mu)$.
-An adjacency matrix $A$ of a graph $\Gamma=(V,E)$ is a $|V|\times|V|$ matrix with rows and columns labelled by the vertices $x,y\in V$, defined by $A_{x,y}=1$ if $x\sim y$ and $A_{x,y}=0$ otherwise.
-Two graphs $\Gamma$ and $\Gamma^{\prime}$ with adjacency matrices $A$ and $A^{\prime}$ are equivalent if there is a permutation matrix $P$ such that $A^{\prime}=P^{T}AP$.
+An adjacency matrix $A$ of a graph $\Gamma=(V,E)$ is a $|V|\times|V|$ matrix with rows and columns labeled by the vertices $x,y\in V$, defined by $A_{x,y}=1$ if $x\sim y$ and $A_{x,y}=0$ otherwise.
+Two graphs $\Gamma$ and $\Gamma^{\prime}$ with adjacency matrices $A$ and $A^{\prime}$ are isomorphic if there is a permutation matrix $P$ such that $A^{\prime}=P^{T}AP$.
+The elementary counting argument gives the following theorem (see [cameron1991designs]).
+Theorem 2.4.
+Let $A$ be an adjacency matrix of $srg(v,k,\lambda,\mu)$. Then $A^{2}=kI+\lambda A+\mu(J-I-A)$, where $I$ an $J$ denote the identity matrix and the all-ones matrix of suitable order, respectively.
 The rank of an adjacency matrix $A$ over $\mathbb{F}_{p}$ is called the $p$-rank of $A$.
@@
 Keys and Rodrigues [Key2018-ec] constructed many LCD codes over $\mathbb{F}_{q}$ from adjacency matrices of strongly regular graphs via orthogonal projectors.
-Especially, they introduced the concept of reflexive LCD codes and complementary LCD codes [Key2018-ec, Definition 2], and provided the condition where adjacency matrices of graphs give LCD codes over $\mathbb{F}_{q}$ [Key2018-ec, Proposition 2].
+In particular, they introduced the concept of reflexive LCD codes and complementary LCD codes [Key2018-ec, Definition 2], and provided the condition where adjacency matrices of graphs give LCD codes over $\mathbb{F}_{q}$ [Key2018-ec, Proposition 2].
+2.3 Two-graphs
+In this subsection, we review the definition and basic properties of two-graphs.
+For a comprehensive treatment on this topic, we refer the reader to [cameron1991designs].
+A two-graph $(V,\Delta)$ consists of a set $V$ whose elements are called vertices, and a collection $\Delta$ of unordered triples of vertices such that each $4$-tuple of $V$ contains an even number of elements of $\Delta$ as subsets.
+Given a simple graph $\Gamma=(V,E)$, the set of triples of $V$ whose induced subgraph in $\Gamma$ contains an odd number of edges gives rise to a two-graph $(V,\Delta)$, and every two-graph can be represented in this way (see [cameron1991designs]).
+Let us introduce another adjacency matrix of a simple graph called “$\mp 1$ adjacency matrix”: A $\mp 1$ adjacency matrix $A$ of a graph $\Gamma=(V,E)$ is a $|V|\times|V|$ matrix with rows and columns labeled by the vertices $x,y\in V$, defined by $A_{x,y}=-1$ if $x\sim y$ and $A_{x,y}=1$ otherwise.
+It is known that simple graphs $\Gamma$ and $\Gamma^{\prime}$, with $\mp 1$ adjacency matrices $A$ and $A^{\prime}$ respectively, represent the same two-graph if there is a diagonal $(1,-1)$-matrix $D$ such that $A^{\prime}=DAD$.
+This defines an equivalence class of graphs, called the switching class, and hence two-graphs can be identified with the switching class of simple graphs.
+The two-graphs $(V,\Delta)$ and $(V^{\prime},\Delta^{\prime})$ are isomorphic if there is a bijection $V\to V^{\prime}$ that induces a bijection $\Delta\to\Delta^{\prime}$.
+Let $\Gamma$ and $\Gamma^{\prime}$ be two simple graphs with $\mp 1$ adjacency matrices $A$ and $A^{\prime}$, respectively.
+By definition, these simple graphs represent isomorphic two-graphs if there are a permutation matrix $P$ and a diagonal $(1,-1)$-matrix such that $A^{\prime}=P^{T}DADP$.
 3 Main results
-In this section, we give main results on a correspondence between binary even LCD codes and adjacency matrices of some simple graphs.
-Also, we provide a lower bound on the minimum weights of binary codes that come from strongly regular graphs.
-3.1 Characterization
+In this section, we establish one-to-one correspondences between binary even LCD codes and certain simple graphs; and ternary LCD codes and certain two-graphs.
+In what follows, we prove general results on orthogonal projectors of LCD codes over $\mathbb{F}_{q}$.
+Specific results on binary and ternary LCD codes will be stated in the following subsections.
 Lemma 3.1.
-Let $C$ be a binary LCD code with orthogonal projector $A$. Then $C$ is the row span of $A$.
+Let $C$ be an LCD code over $\mathbb{F}_{q}$ with orthogonal projector $A$. Then $C$ equals the row span of $A$.
 Proof.
@@
 ∎
-Lemma 3.2.
-Let $C$ and $C^{\prime}$ be binary LCD codes with orthogonal projectors $A$ and $A^{\prime}$, respectively.
-Then, $C$ is equivalent to $C^{\prime}$ if and only if there is a permutation matrix $P$ such that $A^{\prime}=P^{T}AP$.
+Theorem 3.2.
+Let $C$ and $C^{\prime}$ be LCD codes over $\mathbb{F}_{q}$ with orthogonal projectors $A$ and $A^{\prime}$, respectively.
+Then, $C$ is equivalent to $C^{\prime}$ if and only if there is a monomial matrix $N$ such that $A^{\prime}=N^{T}AN$.
 Proof.
 Let $G$ and $G^{\prime}$ be generator matrices of $C$ and $C^{\prime}$ respectively.
-If $C$ and $C^{\prime}$ are equivalent, then there are a monomial matrix $M$ and a permutation matrix $P$ such that $G^{\prime}=MGP$.
-By Theorem 2.1, $A^{\prime}=P^{T}AP$ follows.
-Conversely, suppose that there is a permutation matrix $P$ such that $A^{\prime}=P^{T}AP$.
-Any permutation of the rows of $A$, $P^{T}$ in particular, preserves its row span.
-By Lemma 3.1, the row spans of $A$ and $A^{\prime}$ are $C$ and $C^{\prime}$ respectively.
-It follows that $C^{\prime}=CP$, and hence $C$ and $C^{\prime}$ are equivalent, as required.
+If $C$ and $C^{\prime}$ are equivalent, then there are monomial matrices $M$ and $N$ such that $G^{\prime}=MGN$.
+Now the following computation shows that $A^{\prime}=N^{T}AN$.
+$\begin{split}A^{\prime}&=(MGN)^{T}((MGN)(MGN)^{T})^{-1}(MGN)\\ &=N^{T}G^{T}M^{T}(MGG^{T}M^{T})^{-1}(MGN)\\ &=N^{T}G^{T}(GG^{T})^{-1}GN\\ &=N^{T}AN.\\ \end{split}$
+Conversely, suppose that there is a monomial matrix $N$ such that $A^{\prime}=N^{T}AN$.
+The row span of $A$ is preserved by both permuting coordinates and multiplying any coordinate by nonzero scalar.
+Thus multiplying $A$ by $N^{T}$ from left preserves the row span of $A$.
+Lemma 3.1 tells us that the row spans of $A$ and $A^{\prime}$ are $C$ and $C^{\prime}$, respectively.
+It follows that $C^{\prime}=CN$, and hence $C$ and $C^{\prime}$ are equivalent, as required.
 ∎
-Corollary 3.3.
-Let $C$ and $C^{\prime}$ be binary LCD codes with orthogonal projectors $A$ and $A^{\prime}$, respectively.
-If the traces of $A$ and $A^{\prime}$ are distinct, then $C$ and $C^{\prime}$ are inequivalent.
 Remark 3.3.
-By the proof of Lemma 3.2, it follows that the orthogonal projector of an LCD code is independent of the choice of a generator matrix.
-Remark 3.5.
-Any adjacency matrix $A$ of a simple graph satisfying $A^{2}=A$ is the orthogonal projector of a binary LCD code, where we regard $A$ as a matrix over $\mathbb{F}_{2}$.
-These graphs are characterized by the following conditions:
-  • If two vertices $x$ and $y$ are adjacent, then the number of the common neighbors of $x$ and $y$ is odd.
-  • If two vertices $x$ and $y$ are not adjacent, then the number of the common neighbors of $x$ and $y$ is even.
-Among these graphs there are strongly regular graphs with parameters $(v,k,\lambda,\mu)$, where $k,\mu\equiv 0\pmod{2}$ and $\lambda\equiv 1\pmod{2}$.
+An analogous statement about permutation equivalence was presented in [bardet2019permutation] in the context of public-key cryptography.
 Lemma 3.4.
-Let $C$ be a binary LCD code with orthogonal projector $A$. Then $A$ is an adjacency matrix of a simple graph if and only if $C$ is an even code.
+Let $C$ be an LCD code of length $n$ over $\mathbb{F}_{q}$ with orthogonal projector $A$, and let $r_{i}$ denote the $i$-th row of $A$.
+Then every diagonal entry of $A$ is zero if and only if $(r_{i},r_{i})=0$ for all $1\leq i\leq n$.
 Proof.
-Let $r_{i}$ be the $i$-th row of $A$, and let $n$ be the length of $C$.
-If $A$ is an adjacency matrix of a simple graph, then any diagonal entry of $A$ is zero.
-This fact, along with $A^{2}=A$, implies $(r_{i},r_{i})=0$ for all $1\leq i\leq n$.
-That is, any row of $A$ has an even weight.
-By Lemma 3.1, $C$ is even.
+This is straightforward for the $i$-th diagonal entry of $A^{2}$ equals $(r_{i},r_{i})$. ∎
+3.1 Binary LCD codes and simple graphs
+In this subsection, we restrict ourselves to binary codes. Our aim is to establish the equivalence between binary even LCD codes and certain simple graphs.
+Lemma 3.5.
+Let $C$ be a binary LCD code with orthogonal projector $A$. Then $A$ is an adjacency matrix of some simple graph if and only if $C$ is an even code.
+Proof.
+If $A$ is an adjacency matrix of a simple graph, then any diagonal entry of $A$ is zero. By Lemma 3.4, this implies that any row of $A$ has even weight. Now, the result follows by Lemma 3.1.
 Conversely, suppose $C$ is even.
 Since the orthogonal projector $A$ is a symmetric $(0,1)$-matrix, it remains to show that any diagonal entry of $A$ is 00.
-This can be done by the above argument along with the fact that every codeword of $C$ has an even weight and $A^{2}=A$.
+Since $C$ is even, any codeword of $C$ has even weight.
+In particular, every row of $A$ has even weight, and the result follows by Lemma 3.4.
 This completes the proof.
@@
 Proof.
-This immediately follows from Lemmas 3.1, 3.2 and 3.6. ∎
+The first part of the theorem is an immediate consequence of Lemmas 3.1 and 3.5.
+To prove the second part, recall that simple graphs $\Gamma$ and $\Gamma^{\prime}$ with adjacency matrices $A$ and $A^{\prime}$ respectively are isomorphic if there is a permutation matrix $P$ such that $A^{\prime}=P^{T}AP$.
+Since any permutation matrix is a monomial matrix, Theorem 3.2 implies the result.
+∎
+Any adjacency matrix $A$ of a simple graph satisfying $A^{2}=A$ is the orthogonal projector of a binary LCD code, where we regard $A$ as a matrix over $\mathbb{F}_{2}$.
+These graphs are characterized by the following conditions:
+  • If two vertices $x$ and $y$ are adjacent, then the number of the common neighbors of $x$ and $y$ is odd.
+  • If two vertices $x$ and $y$ are not adjacent, then the number of the common neighbors of $x$ and $y$ is even.
+Among these graphs there are strongly regular graphs with parameters $(v,k,\lambda,\mu)$, where $k,\mu\equiv 0\pmod{2}$ and $\lambda\equiv 1\pmod{2}$.
 Remark 3.7.
@@
 Hence it follows that $(A)_{i,i}=0$ if and only if $C^{i}$ is LCD, and $A_{i,i}=1$ if only if $C_{i}$ is LCD.
-3.2 Bounds on minimum weights
-In this subsection, we give a lower bound on the minimum weights of binary codes obtained as row spans of adjacency matrices of strongly regular graphs.
+3.2 Ternary LCD codes and two-graphs
+In this subsection, we restrict ourselves to ternary codes.
+Our aim is to establish the equivalence between ternary LCD codes and certain two-graphs.
+Throughout this subsection, the term “adjacency matrix” means the $\mp 1$ adjacency matrix (see Section 2).
+Lemma 3.8.
+Let $C$ be a ternary LCD code with orthogonal projector $A$. Then $A$ is an adjacency matrix of some two-graph if and only if every row of $A$ has weight divisible by three.
+Proof.
+If $A$ is an adjacency matrix of a two-graph, then any diagonal entry of $A$ is zero.
+The result follows by Lemma 3.4 and the fact that $(x,x)=0$ if and only if $\operatorname{wt}(x)\equiv 0\pmod{3}$ for any $x\in\mathbb{F}_{3}^{n}$.
+Conversely, suppose that any row of $A$ has weight divisible by three.
+Since the orthogonal projector $A$ is a symmetric matrix whose entry is in $\{0,1,-1\}$, it remains to show that any diagonal entry of $A$ is 00.
+Again, the result follows by Lemma 3.4 and the fact that $(x,x)=0$ if and only if $\operatorname{wt}(x)\equiv 0\pmod{3}$ for any $x\in\mathbb{F}_{3}^{n}$.
+This completes the proof.
+∎
+In both binary and ternary cases, the conditions for orthogonal projectors to be adjacency matrices are characterized by weights of their row vectors.
+In binary case, Lemma 3.1 implies every codeword to have even weight, while an analogous statement is false in ternary case.
+In fact, if every codeword in $C$ has weight divisible by three, then $C$ is self-dual code and hence cannot be LCD.
+The condition “every row of the orthogonal projector has weight divisible by three” is not insightful since such a statement is seemingly not related to any property of the corresponding code.
+From this perspective, the following sufficient condition is more useful.
+Theorem 3.9.
+Let $C$ be a ternary LCD code of length $n$ with orthogonal projector $A$. If $A$ is an adjacency matrix of some two-graph, then the dimension of $C$ is divisible by three.
+Proof.
+In this proof, we regard $A$ as a matrix over $\mathbb{F}_{3}$.
+Since $A^{2}=A$, the eigenvalue of $A$ is either 00 or $1$.
+To see this, let $a$ be any eigenvalue of $A$ and $v$ be an eigenvector with respect to $a$.
+Then $0=(A^{2}-A)v=(a^{2}-a)v$, and hence $a\in\{0,1\}$, as required.
+Therefore, there is an invertible matrix $P$ over $\mathbb{F}_{3}$ such that $P^{-1}AP=\operatorname{diag}{(1,\dots,1,0,\dots,0)}$.
+Let $D$ denote this diagonal matrix.
+The code $C$ is exactly the $A$-invariant subspace, so its dimension equals the number of $1$s in $D$, which we denote by $k$.
+Now, the proof is concluded by the facts $\operatorname{tr}(A)=\operatorname{tr}(D)$ and $\operatorname{tr}(D)\equiv k\pmod{3}$, where in the latter we regard $\operatorname{tr}(D)$ as an integer.
+∎
+The converse of the above theorem is false: Let $C$ be a ternary code whose generator matrix is of the form $G=\begin{pmatrix}I_{k}&A\end{pmatrix},$ where $k\equiv 0\pmod{3}$ and $A$ is $k\times(n-k)$ matrix with $AA^{T}=O$.
+Since $GG^{T}=I_{k}$ is invertible, $C$ is LCD, and one can easily check that
+$G^{T}(GG^{T})^{-1}G=\begin{pmatrix}I_{k}&A\\ A^{T}&O\end{pmatrix}.$
+Because of the k $1$s in its diagonal, this matrix fails to be an adjacency matrix of any two-graph.
+Theorem 3.10.
+The orthogonal projector $A$ of a ternary LCD code is an adjacency matrix of a simple graph satisfying $A^{2}=A$, where we regard $A$ as a matrix over $\mathbb{F}_{2}$.
+Conversely, such an adjacency matrix is the orthogonal projector of a ternary LCD code.
+Furthermore, two binary even LCD codes are equivalent if and only if corresponding graphs are isomorphic.
+Proof.
+The first part of the theorem is an immediate consequence of Lemmas 3.1 and 3.5.
+To prove the second part, let $\Gamma$ and $\Gamma^{\prime}$ be two-graphs with adjacency matrices $A$ and $A^{\prime}$, and suppose their rows generate ternary LCD codes $C$ and $C^{\prime}$, respectively.
+If $\Gamma$ and $\Gamma^{\prime}$ are isomorphic, then there are a diagonal $(1,-1)$-matrix $D$ and a permutation matrix $P$ such that $A^{\prime}=P^{T}DADP$.
+If we write $N=DP$, then $N$ is a monomial matrix and $A^{\prime}=N^{T}AN$.
+Hence $C$ and $C^{\prime}$ are equivalent by Theorem 3.2.
+Conversely, if $C$ and $C^{\prime}$ are equivalent, then there is a monomial matrix $N$ such that $A^{\prime}=N^{T}AN$, again by Theorem 3.2.
+Recall that any monomial matrix can be represented by a diagonal matrix and a permutation matrix.
+Hence there are a diagonal $(1,-1)$-matrix $D$ and a permutation matrix $P$ such that $N=DP$.
+Therefore, we have $A^{\prime}=P^{T}DADP$, as required.
+∎
+4 Bounds on minimum weights
+In this section, we give a lower bound on the minimum weights of binary codes obtained as row spans of adjacency matrices of strongly regular graphs.
 These results will be applied in Section 5.
-Note that all the assertions in this subsection hold not only for binary LCD codes but also for any binary codes.
+Note that all assertions in this subsection hold not only for binary LCD codes but also for any binary codes.
 Proposition 4.1.
@@
 5 Applications
-In this section, we illustrate examples where results from the preceding section can be applied.
-4.1 Binary codes from the Paley graph
-Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of adjacency matrices of strongly regular graphs are characterized by their eigenvalues.
-For strongly regular graphs $srg(41,20,9,10)$, any corresponding binary code has $2$-rank, namely, dimension $20$.
-Haemers, Peeters and Rijckevorsel [Haemers1999-no] considered binary codes obtained as row spans of strongly regular graphs.
-They observed that for $srg(25,12,5,6)$ and $srg(41,20,9,10)$, any pair of non-isomorphic graphs gave a pair of inequivalent codes, and among such codes the code from the Paley graph had the largest minimum weight.
-Based on this result, they conjectured that the latter statement holds in general (see [Haemers1999-no, Section 4.3]).
-The classification of $srg(25,12,5,6)$ has been completed [paulus1973conference]; there are exactly $15$ non-isomorphic $srg(25,12,5,6)$, and with the aid of computer it was shown in [Haemers1999-no, Section 6.1] that the Paley graph gives the code with largest minimum weights.
-On the other hand, the classification of $srg(41,20,9,10)$ has not been completed yet.
-However, it is not difficult to show the Paley graph of order $41$ gives a binary even $[41,20]$ code whose minimum weight is largest among all $srg(41,20,9,10)$.
-One way to see this is to use the Griesmer bound, which tells us the minimum weight of binary LCD $[41,20]$ code is at most $11$.
-By Lemma 3.6, the minimum weight must be even, and hence the result follows.
-According to [grassl_bounds_nodate], the largest minimum weight of binary linear code is exactly $10$, so it is obvious that the code from the Paley graph of order $41$ attains the largest minimum weight.
-To best of our knowledge, the conjecture is still open for $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ with $q>41$ and $q\equiv 1\pmod{8}$.
-Remarkably, the first part of the above observation by Haemers, Peeters and Rijckevorsel holds in general for any strongly regular graph with parameters $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ for $q\equiv 1\pmod{8}$ since any adjacency matrix of such a graph satisfies the condition given in Theorem 3.7.
-We state this fact as the following corollary:
+Haemers, Peeters, and Rijckevorsel [Haemers1999-no] considered binary codes obtained as row spans of strongly regular graphs. With the help of computers, they observed that
+  1. (i)
+    for $srg(25,12,5,6)$ and $srg(41,20,9,10)$, any pair of non-isomorphic graphs gave a pair of inequivalent codes;
+  2. (ii)
+    among such codes the code from the Paley graph had the largest minimum weight.
+Based on these results, they conjectured that the latter statement holds in general (see [Haemers1999-no, Section 4.3]).
+In this section, we (i) prove that the former statement is true for any strongly regular graph with parameters $(q,(q-1)/2,(q-5)/4,(q-1)/4)$ for $q\equiv 1\pmod{8}$, and (ii) verify the latter statement for $srg(41,20,9,10)$.
+Note that the result by Brouwer and Eijl [Brouwer1992-nc] shows that for strongly regular graphs $srg(25,12,5,6)$ and $srg(41,20,9,10)$, corresponding binary codes have dimension $12$ and $20$, respectively.
+First, we verify that the first part of the above observation by Haemers, Peeters and Rijckevorsel holds in general for any strongly regular graph with parameters $(q,(q-1)/2,(q-5)/4,(q-1)/4)$ for $q\equiv 1\pmod{8}$.
+This can be done easily using the main results of this paper.
 Corollary 5.1.
 Any pair of non-isomorphic $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ with $q\equiv 1\pmod{8}$ gives a pair of inequivalent binary even LCD codes, and vice versa.
+Proof.
+Let $A$ be an adjacency matrix of $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ with $q\equiv 1\pmod{8}$.
+By Theorem 2.4 we have $A^{2}=\frac{q-1}{2}I+\frac{q-5}{4}A+\frac{q-1}{4}(J-I-A)$, where $I$ and $J$ denote the identity matrix and the all-ones matrix of suitable order, respectively.
+Since we assume that $q\equiv 1\pmod{8}$, $A^{2}=A$ holds when we regard $A$ as a matrix over $\mathbb{F}_{2}$.
+This fact, together with Theorem 3.6, implies that the resulting code is an even LCD code.
+Hence the result follows.
+∎
+Next, we study the second part of the observation above.
+We remark that the classification of $srg(25,12,5,6)$ has been completed [paulus1973conference]; there are exactly $15$ non-isomorphic $srg(25,12,5,6)$, and with the aid of computer it was shown in [Haemers1999-no, Section 6.1] that the Paley graph gives the code with largest minimum weights.
+On the other hand, the classification of $srg(41,20,9,10)$ has not been completed yet, making it difficult to prove the problem in a similar way.
+Theorem 5.2.
+The Paley graph of order $41$ gives a binary even $[41,20]$ code whose minimum weight is largest among all $srg(41,20,9,10)$.
+Proof.
+The result by Brouwer and Eijl [Brouwer1992-nc] shows that for $srg(41,20,9,10)$ any corresponding binary code has dimension $20$, respectively, and the Griesmer bound tells us the minimum weight of binary LCD $[41,20]$ code is at most $11$.
+Recall that by the proof of Corollary 5.1, any $srg(41,20,9,10)$ gives an even LCD code, and hence the minimum weight must be at most $10$.
+A computer search verifies that the Paley graph gives a code which attains this bound.
+This concludes the proof.
+∎
+According to [grassl_bounds_nodate], the largest minimum weight of binary linear code is exactly $10$, so we can simply use this fact to prove Theorem 5.2.
+However, the proof of this bound depends on a lengthy computer search, making it difficult to understand.
+From this perspective, the proof presented in this paper is more sophisticated.
+To the best of our knowledge, the conjecture is still open for $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ with $q>41$ and $q\equiv 1\pmod{8}$.
+Remark 5.3.
+It is unknown whether there is any other $srg(41,20,9,10)$ than the Paley graph which attains the bound. The author thinks that this problem is interesting.
 According to [Maksimovic2022-xa], there are at least $7152$ strongly regular graphs with parameters $(41,20,9,10)$ up to isomorphism.
-We constructed binary even LCD codes from the $7152$ graphs, and verified the following:
-  • There are mutually inequivalent,
-  • Only the code that comes from the Paley graph has the largest minimum weight $10$,
-  • All codes have minimum weight greater than $4$, and the code comes from the following adjacency matrix has the minimum weight $4$.
-By Proposition 3.10, any code from $srg(41,20,9,10)$ must have the minimum weight larger than $d\geq 1+(k+1)/(\min\{\lambda,\mu\}+1)=3$.
+We constructed binary even LCD codes from the $7152$ graphs, and verified that (i) there are mutually inequivalent; (ii) only the code that comes from the Paley graph has the largest minimum weight $10$; and (iii) all codes have minimum weight greater than $4$, and the code comes from the following adjacency matrix has the minimum weight $4$.
+By Proposition 4.2, any code from $srg(41,20,9,10)$ must have minimum weight larger than $d\geq 1+(k+1)/(\min\{\lambda,\mu\}+1)=3$.
 Applying Theorem 3.6, we can improve the lower bound by $1$, and hence the lower bound on the minimum weight is $4$.
 The above observation shows this lower bound is tight for $srg(41,20,9,10)$.
-4.2 Optimal LCD codes from regular graphs
-In this subsection, we improve some lower bounds on the minimum weight of binary LCD codes by constructing optimal binary LCD codes from their orthogonal projectors.
-For recent results on optimal binary LCD codes, we refer the reader to [bouyuklieva_optimal_2021].
-Let $d(n,k)$ denote the largest minimum weight among all LCD $[n,k]$ code.
-Here we consider a construction by circulant matrices. A circulant matrix $C$ of order $n$ is a matrix of the following form
-$C=\begin{pmatrix}c_{0}&c_{1}&c_{2}&\cdots&c_{n-2}&c_{n-1}\\ c_{n-1}&c_{0}&c_{1}&\cdots&c_{n-3}&c_{n-2}\\ c_{n-2}&c_{n-1}&c_{0}&\cdots&c_{n-4}&c_{n-3}\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ c_{2}&c_{3}&c_{4}&\cdots&c_{0}&c_{1}\\ c_{1}&c_{2}&c_{3}&\cdots&c_{n-1}&c_{0}\\ \end{pmatrix}.$
-Circulant matrices are determined by their first rows, and hence there are $2^{n}$ circulant matrices over $\mathbb{F}_{2}$ of order $n$.
-For $26\leq n\leq 37$, we constructed all circulant matrices over $\mathbb{F}_{2}$ of order $n$, which must be checked further for the condition given in Theorem 3.7.
-Then by Lemma 3.1 we constructed the binary LCD codes from the orthogonal projectors, and computed their minimum weights.
-Consequently, we obtain the following theorem:
-Theorem 4.2.
-There are optimal binary LCD $[n,k,d]$ codes for
-$(n,k,d)=$ $(31,10,10),(31,11,10),(31,20,6),(33,10,12),$
-$(33,11,11),(33,12,10),(33,13,10),(33,20,6),(33,22,6).$
-For each of these codes, the first rows of their orthogonal projectors are listed in Table 1.
-For a binary LCD $[n,k,d]$ code $C$, it is easy to see that $C^{\prime}=\{(0,c)\mid c\in C\}$ is a binary LCD $[n+1,k,d]$ code.
-Using this property, we obtained three optimal binary LCD codes from some codes listed in Theorem 4.2.
-Consequently, we obtain the following theorem:
-Theorem 4.3.
-There are optimal binary LCD $[n,k,d]$ codes for
-$(n,k,d)=(33,21,6),(34,13,10),(34,22,6).$
-Any orthogonal projector mentioned above has constant diagonal entries, and hence any binary LCD code in Theorem 4.2 is either an even code or its dual code.
-Note that the orthogonal projectors of LCD codes above (or their dual codes) are adjacent matrices of regular graphs by construction.
-It is remarkable that adjacency matrices of regular graphs can give LCD codes with the largest minimum weights.
-Furthermore, we remark that any code in Theorem 4.2 has odd length.
-This follows by the fact that the dimension of any binary even LCD code is even, and that the dual code of any even code must contain the all-ones vector; if a binary even LCD $[n,k]$ code $C$ has even length, then the all-ones vector must be in $C\cap C^{\perp}$, a contradiction.
-Table 1: Optimal binary even LCD codes
-$n$ $k$ $d$ the first row
-$31$ $10$ $10$ $(0000010001110100001011100010000)$
-$31$ $11$ $10$ $(1001011101111110011111101110100)$
-$31$ $20$ $6$ $(0001011101111110011111101110100)$
-$33$ $10$ $12$ $(000101110111111100111111101110100)$
-$33$ $11$ $11$ $(100001010010011000011001001010000)$
-$33$ $12$ $10$ $(000101110110111100111101101110100)$
-$33$ $13$ $10$ $(100001010011011000011011001010000)$
-$33$ $20$ $6$ $(000001010011011000011011001010000)$
-$33$ $22$ $6$ $(000001010010011000011001001010000)$
 6 Concluding remarks
-In this paper, we proved that binary even LCD codes and some graphs are in one-to-one correspondence in a certain way.
-Furthermore, we provided lower bounds on minimum weights of codes obtained from strongly regular graphs.
-Consequently, we gave an explanation to the observation by Haemers, Peeters, Rijckevorsel [Haemers1999-no].
-Our result can be applied to generate a graph that gives optimal binary code, as shown in Section 4.
-We hope that studying such graphs leads to a discovery of a further relation between graphs and even LCD codes.
+In this paper, we have established one-to-one correspondences between binary even LCD codes and certain simple graphs; and ternary LCD codes and certain two-graphs.
+Based on these results, we further studied the relation between LCD codes and graphs.
+Our motivation is to study whether a good graph gives a rise to a good code.
+We verified that for some strongly regular graphs the answer is positive, for (i) the automorphism groups of graphs equal those of the resulting codes, and (ii) the Paley graph of order $41$ gives a binary even $[41,20]$ code whose minimum weight is largest among all codes from $srg(41,20,9,10)$.
+It is unknown whether there is any other $srg(41,20,9,10)$ than the Paley graph which attains the bound, and the author thinks that this problem is interesting.
 7 Acknowledgement
-The author would like to thank Stefka Bouyuklieva for several questions and comments that improved the manuscript.
+The author would like to thank Stefka Bouyuklieva for her insightful questions and comments, which improved the manuscript.